OpenStudy (anonymous):

find the slope of the curve y=x^2+x at x=3

OpenStudy (anonymous):

if you could assist me i would greatly appreciate your effort.@amistre64 @ganeshie8 @myininaya @jim_thompson5910

OpenStudy (anonymous):

@amistre64

OpenStudy (tkhunny):

Have you considered the 1st Derivative?

OpenStudy (anonymous):

yes i beleive im doing it wrong as i get (3+h)^2+3-12/h and end up getting the slope to be six but thats not one of my choices

OpenStudy (anonymous):

slope is nothing but derivative that is dy/dx so can you tell me what is derivative of x^2 + x ?

OpenStudy (anonymous):

im not familiar with that formula as i have just begun calculus and we have only been taught the difference quotient of f at a. @tgawade

OpenStudy (anonymous):

use this derivative of x^n = n * x^(n-1)

OpenStudy (anonymous):

so can u tell me what is derivative of x^2 = --?

hartnn (hartnn):

[(3+h)^2+3-12]/h is correct, how did you proceed ?

OpenStudy (anonymous):

i simplified and got 6h+h^2/h

OpenStudy (anonymous):

then i cancel the hs and got 6+h

OpenStudy (anonymous):

then i pugged in zero and got 6

hartnn (hartnn):

let me go through it again

OpenStudy (anonymous):

should i have replaced both x's in x^2+x with 3+h instead of one

hartnn (hartnn):

f(x+h) = (x+h)^2 +(x+h) so [(3+h)^2+3-12]/h is infact NOT correct

hartnn (hartnn):

yes, thats correct, both "x"s in f(x), replace it by x+h

hartnn (hartnn):

[(3+h)^2+3+h-12]/h

OpenStudy (anonymous):

so you would get h(6+h+1)/h cancel hs then plug in o for h and get 7 for the slope correct

hartnn (hartnn):

now you will get the correct answer :) oh you already did! yesss, 7 is correct slope :)

OpenStudy (anonymous):

thank you for your assistance i appreciate your time and effort @hartnn @tgawade @tkhunny

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

i know the easiest way to solve this does anybody want to know ??????